# A New Perspective on Double-S Curve Motions of Higher Order and Optimal Motion Planning

### v2

@@ Abstract
This paper presents and proves an equation for the time horizon of symmetric trajectories with zero boundary conditions and bounded derivatives of arbitrary order. This equation holds regardless of the number of phases comprising the associated motion. This avoids case distinctions in calculations. Application examples of motions with minimum time, minimum velocity, and minimum acceleration are discussed. Furthermore, an algorithm is derived that reduces the time minimization problem to solving a system of equations. This algorithm avoids nested case distinctions and complex optimizations.
@@ 1 Introduction
Optimal motion planning has been done for many years, and it is still an active research field – particularly from the mathematical-algorithmical and the numerical-computational perspective. Some important applications are robotics (see [1, 2]), aeronautics and astronautics (see [3, 4]), or transport planning (see [5]) – just to mention a few. A rather simple but important task is to find the minimum-time motion for accomplishing a given distance $s$ under certain restrictions: e.g., the velocity is bounded, or the acceleration is bounded, or both are (see [6] for an excellent book about the kinematics of trajectory planning). Common approaches consider the jerk as third derivative as well [7, 8, 9]. Such approaches work in a 7-segment motion (SSM) profile (which is also referred as double-S curve, see [10] for early work and [11, 12] as well) to find relations between the minimal time, the distance and the boundary values for several cases and under certain boundary conditions. Fourth order motions are considered in [13, 14].
The aim of this paper is to reveal new perspectives on symmetric [^1] For a treatment of asymmetric trajectories see [15]. double-S curves of arbitrary order. Following the work of [16] (see also [17, 18]), a general equation for the duration of the motion $T$ is proven. In the case of jerk-controlled motion [^2] The term “jerk-controlled” means that the velocity, the acceleration and the jerk are assumed to be bounded., this equation is
@@
Moreover, since the equation is true for all symmetric trajectories with vanishing boundary values, it can be used to describe not only the time-minimal motion over the distance $s$. Conversely, if a time horizon and a distance are given, the question arises as to which trajectory – among all possible motions – minimizes, for example, the achieved velocity (or acceleration, or jerk, etc.). Equation (1) and its generalization are also suitable for such questions. In the following, a systematic approach to various optimizations and arbitrary orders is developed using a single equation.
@@ 2 Problem Setup
Let us consider one-dimensional trajectories encoded in $x(t)$ where $x\in\mathbb{R}$ denotes the position as a function of the time $t$ which ranges in the interval $[0,T]$ with $T>0$. In addition, the order parameter is denoted by $N$ with $N\in\mathbb{N}$ and $N>0$. We assume $x(t)$ to be (N1)(N-1)-times continuously differentiable w.r.t. $t$, and the $N$-th derivative of $x(t)$ should be piecewise continuous with countable many finite jumps. Then, $\dot{x}(t)$ is the velocity, and – if existing – $\ddot{x}(t)$ and $\dddot{x}(t)$ stand for the acceleration and the jerk, respectively [^4] The dot indicates derivatives w.r.t. $t$. And in general, the $k$-th derivative w.r.t. $t$ is written as $x^{(k)}(t).$.
@@
For both MTT and MDT the key characteristics of the optima are known [9]. Basically, the $N$-th derivative of an optimal trajectory is piecewise constant [^6] This is due to Pontryagin’s maximum principle, see [19] for instance., and optimal trajectories enjoy a certain kind of symmetry in such a way that – for example – accelerating and decelerating are time-reserved processes. The following definition completes this section by formalizing the above mentioned relevant trajectory type.
@@
Obviously, the constant $c$ in Definition 2 (iv) should be less or equal than the bound $w_{N}$ (see Definition 1 (iv)).
The next section recalls briefly the pre-optimal trajectories up to the 7-segment motion at an elementary level.
3 Seven-Segment Motions
Before we tackle the general proof, we briefly recall pre-optimal trajectories within the SSM framework while climbing from $N=1$ to $N=3$.
@@ 3.1 Acceleration-Controlled Motions
Let $v$ denote the maximum velocity reached during the motion, i.e. $v=\max\limits_{t\in[0,T]}|\dot{x}(t)|$. In the triangular case we have
@@ 3.2 Jerk-Controlled Motions
By including the jerk constraint $(N=3)$ we now reach the actual case of the SSM. However, there are four cases to be distinguished (see Figure 2). We derive Equation (1) for each of the cases. Let the pre-optimal constant $c$ be denoted by $j$, let $v$ be defined as before in Section 3.1, and let $a:=\max\limits_{t\in[0,T]}|\ddot{x}(t)|$ denote the maximal acceleration reached during the motion.
@@
4-Phase Case
Clearly,
@@
$x(t)$ $=\frac{1}{6}jt_{1}^{3}+\frac{1}{2}v(t-t_{1})+\frac{1}{2}a(t-t_{1})^{2}-\frac{1}{6}(t-t_{1})^{3}$ (11)
@@
5-Phase Case
Again Equation (7) is valid, but now we have with Equation (11) that
@@
6-Phase Case
In this case,
@@
and via Equation (15) to $\frac{s}{v}=\frac{T}{2}$. Applying Equation (23) leads to $\frac{v}{a}+\frac{a}{j}=\frac{T}{2}$ as well, which again proves Equation (1).
7-Phase Case
From the 6-phase case we can already adapt
@@
which completes the proof of Equation (1) for all cases of SSM.
However, since scanning larger decision trees seems quite unsatisfactory, the next section presents a systematic approach for arbitrary orders.
@@ 4 Time Horizon of Pre-Optimal Motions
This section presents and proves the main result:
@@
With the notation from Theorem 3, we introduce the earlest time point of achievement as
$T_{n}=\min\left\{t\in[0,T]|x^{(n)}(t)=x_{n}\right\}$ (31)
@@
$T_{n}$ $=$ $\min\left\{t\in[0,T]|x^{(n)}(t)=x_{n}\right\}$ (42)
$=$ $\min\left\{t\in[0,T]|f_{N-n}(t)=f_{N-n}(t_{N-n-1})\right\}$
@@ 5 Applications and Examples
Equipped with the general Equation (32) of Theorem 5 we are able to gain some new perspectives of optimal motion planning. In the first subsection we will briefly discuss the cases $N=1$ (velocity-controlled motions) and $N=2$ (acceleration-controlled motions) before the power and elegance of Equation (1) are fully exploited while treating jerk-controlled and snap-controlled movements (Subsections 5.2 and 5.3, respectively). Section 5 closes with remarks on arbitrary order (Subsection 5.4).
@@ 5.1 Velocity-Controlled and Acceleration-Controlled Motions
  • $\hat{v}_{\mathrm{MTT}}=\hat{v}^{\star}$, $\hat{T}_{\mathrm{MTT}}=\hat{T}^{\star}$ if $\hat{v}^{\star}\leq\hat{v}_{\max}$ (triangular case in Section 3.1, see Figure 1 left), or
  • $\hat{v}_{\mathrm{MTT}}=\hat{v}_{\max}$, $\hat{T}_{\mathrm{MTT}}=\frac{1}{\hat{v}_{\max}}+\hat{v}_{\max}$ otherwise (trapezoidal case, see Figure 1 right).
Figure 3: $\hat{T}$ as a function of $\hat{v}$ (black curve) according to Equation (48) which exhibits a global minimum at $\hat{v}^{\star}=1$ and $\hat{T}^{\star}=2$ (blue cross). The triangular case of MTT occurs if $\hat{v}_{\max}$ lies behind this minimum (green dashed line). Then the minimal time is exactly $\hat{T}^{\star}$. Otherwise (magenta dashed line, trapezoidal case) the optimum of MTT is located at the magenta bullet. The MDT optimization is colored in orange. The optimum is indicated by the orange bullet.
@@ 5.2 Jerk-Controlled Motions
The “fish diagram” of Figure 4 displays $\hat{T}$ as a function of $\hat{v}$ and $\hat{a}$ as a contour plot. There are two red curves – $\hat{a}_{\min}$ curve and $\hat{v}_{\min}$ curve – which frame the region of all principally possible motions. These curves are defined by the conditions (see Equation (8)):
@@
If such a condition is satisfied the associated order has no cruising phase. For example, $\hat{T}_{1}=2\hat{T}_{2}$ means that $\ddot{x}$ has a triangular shape if it is not equal to 0 (see Figure 2 both upper panels). If there is no cruising phase in $\ddot{x}$, the time for reaching $\hat{v}$ (equal to $\hat{T}_{1}$) is exactly twice the time needed for reaching $\hat{a}$ (equal to $\hat{T}_{2}$). This corresponds to the fact that for every acceleration a certain velocity is quasi automatically achieved. Therefore, not all of the ($\hat{v},\hat{a}$)-pairs are actually possible. And it is precisely these borderline cases that are reflected in the conditions (58, 59).
Furthermore, the global minimum is determined by simultaneously satisfying the equations (58, 59). It is located at $\hat{a}^{\star}=\sqrt[3]{\frac{1}{2}}$, $\hat{v}^{\star}=\sqrt[3]{\frac{1}{4}}$ and $\hat{T}^{\star}=\sqrt[3]{32}$.
Let us now take a closer look at MTT. The four possible cases discussed in Section 3.2 can be easily recovered and reinterpreted within the fish diagram. As above, $\hat{v}_{\max}$ and $\hat{a}_{\max}$ are horizontal and vertical cut-offs, respectively. Then:
@@
There is a similar procedure for MDT. Since $\hat{T}_{\mathrm{MDT}}$ is given we work on a fixed contour curve.
@@
All in all, the relevant optimal values can be obtained by targeted solving of algebraic equations. In addition, the fish diagram provides an overview of which case actually occurs, i.e. of how many phases the motion consists and which particular bounds are active. Noteworthy, independent of the cases and MTT or MDT, Equation (51) remains overall valid.
@@
Figure 6: Schematic illustration of the solution of MDT with $M=1$, i.e. minimum-velocity motions. In any case, the solution is indicated by an orange bullet. Left: if $\hat{a}_{\max}>\hat{a}^{\star}$, the minimal velocity is achieved at the point where the contour curve belonging to the given time horizon $\hat{T}_{\mathrm{MDT}}$ leaves the admissible region and intersects the $\hat{v}_{\max}$ curve. Right: For small enough values of $\hat{a}_{\max}$ the contour curve reaches the acceleration boundary before the $\hat{v}_{\min}$ are. The solution is determined by this intersection.

### v3

@@ Abstract
This paper presents and proves an equation for the time horizon of symmetric trajectories with zero boundary conditions and bounded derivatives of arbitrary order. This equation holds regardless of the number of phases comprising the associated motion, which avoids case distinctions in calculations. Application examples of motions with minimum time, minimum velocity, and minimum acceleration are discussed. Furthermore, an algorithm is derived that reduces the time minimization problem to solving a system of equations. This algorithm avoids nested case distinctions and complex optimizations.
@@ 1 Introduction
Optimal motion planning has been done for many years, and it is still an active research field – particularly from the mathematical-algorithmical and the numerical-computational perspective. Some important applications are robotics (see [1, 2]), aeronautics and astronautics (see [3, 4]), or transport planning (see [5]) – just to mention a few. A rather simple but important task is to find the minimum-time motion for accomplishing a given distance $s$ under certain restrictions: e.g., the velocity is bounded, or the acceleration is bounded, or both (see [6] for an excellent book about the kinematics of trajectory planning); surveys in the field of optimization-based motion planning are, e.g., [7, 8]. Common approaches consider the jerk as third derivative as well [9, 10, 11]. Such approaches work in a 7-phase motion (SPM) profile (which is also referred as 7-phase motion or double-S curve, see [12] for an early work and [13, 14] as well) to find relations between the minimal time, the distance and the boundary values for several cases and under certain boundary conditions. Fourth order motions, i.e. snap-controlled motions with constraints for the first four derivations of the position, are considered in [15, 16]. Particularly, double-S (or simply S-curve) motion planning remains a vital topic in high-precision robotic trajectory generation: for instance [17] propose a time-optimal asymmetric S-curve trajectory for redundant manipulators under kinematic constraints, using a whale-optimization algorithm to schedule jerk-limited motion phases while respecting both end-effector and joint limits. In parallel, [18] introduce a general asymmetric S-curve profile optimized to minimize execution time while keeping residual vibration low, demonstrating analytic performance gains over symmetric or simpler S-curve profiles. Moreover, [17] present a time-optimal, jerk-continuous trajectory planning method for robot manipulators that enforces third-order kinematic constraints (velocity, acceleration, jerk) to produce smooth, practically trackable motions. In industrial motion-planning scenarios, [19] propose an SPM optimizer under joint kinematics limits and waypoint constraints, framing the trajectory planning as a nonlinear optimization to minimize cycle time. Recent work also pushes beyond purely time-optimal planning: [20] develop a multi-objective optimal trajectory planner for manipulators that jointly optimizes time and jerk (or smoothness), using quintic B-splines and a meta-heuristic to respect kinematic constraints while trading off between speed and smoothness.
The aim of this paper is to reveal new perspectives on symmetric double-S curves of arbitrary order. For a treatment of asymmetric trajectories see [21]. Following the work of [22] (see also [23, 24]), a general equation for the duration of the motion $T$ is proven. In the case of jerk-controlled motion [^1] The term “jerk-controlled” means that the velocity, the acceleration and the jerk are assumed to be bounded., this equation is
@@
Moreover, since the equation is true for all symmetric [^3] I.e. roughly speaking acceleration = deceleration, see below. trajectories with vanishing boundary values, it can be used to describe not only the time-minimal motion over the distance $s$. Conversely, if a time horizon and a distance are given, the question arises as to which trajectory – among all possible motions – minimizes, for example, the achieved velocity (or acceleration, or jerk, etc.). Equation (1) and its generalization are also suitable for such questions. In the following, a systematic approach to various optimizations and arbitrary orders is developed using a single equation.
@@ 2 Problem Setup
Let us consider one-dimensional trajectories encoded in $x(t)$ where $x\in\mathbb{R}$ denotes the position as a function of the time $t$ which ranges in the interval $[0,T]$ with $T>0$. In addition, the order parameter is denoted by $N$ with $N\in\mathbb{N}$ and $N>0$. We assume $x(t)$ to be (N1)(N-1)-times continuously differentiable w.r.t. $t$, and the $N$-th derivative of $x(t)$ should be piecewise continuous with countable many finite jumps. Then, $\dot{x}(t)$ is the velocity, and – if existing – $\ddot{x}(t)$ and $\dddot{x}(t)$ stand for the acceleration and the jerk, respectively [^4] According to Newton, the dot indicates derivatives w.r.t. $t$. And in general, the $k$-th derivative w.r.t. $t$ is written as $x^{(k)}(t).$.
@@
In what follows, we have to distinguish between the given bounds $w_{n}$ which are sometimes more concretely denoted by $v_{\max},a_{\max},j_{\max}$ etc. and actually achieved values $\max\limits_{t\in[0,T]}\left|x^{(n)}(t)\right|$. These values will be denoted by $v,a,j$ etc. for $n=1,2,3$, respectively. Some authors call the letters ones “peak values” (in contrast to “maximum values” for $w_{n}$) which may be quite misleading if – for example – the velocity $v$ is held for a certain time interval (cruising phase).
@@
For both MTT and MDT the key characteristics of the optima are known [11]. Basically, the $N$-th derivative of an optimal trajectory is piecewise constant [^6] This is due to Pontryagin’s maximum principle, see [25] for instance., and optimal trajectories enjoy a certain kind of symmetry in such a way that – for example – accelerating and decelerating are time-reversed processes. The following definition completes this section by formalizing the above mentioned relevant trajectory type.
@@
A supporting interval is an interval, where the respective derivative does not vanish. If a supporting interval is maximal it can not be extended to the left or the right. That means that the boundary values of the respective derivative are zero for a maximal supporting interval. In Figure 1, right panel, any subinterval of $[0,T]$ is a supporting interval of $\dot{x}$ but only $[0,T]$ is a maximal supporting interval. Regarding the acceleration $\ddot{x}$, the maximal supporting intervals are $0,t_{1}]$ and $T-t_{1},T]$. The proposed symmetry of Definition 2 3 can be seen from that panel as well for $\dot{x}$. This symmetry means that any order restricted to a maximal supporting interval is axisymmetric to a vertical axis drawn through the middle of the interval. Definition 2 (iv) means that the highest order is piecewise constant. Obviously, the constant $c$ in Definition 2 (iv) should be less than or equal to the bound $w_{N}$ (see Definition 1 (iv)).
The next section recalls briefly the pre-optimal trajectories up to the 7-phase motion at an elementary level.
3 Seven-Phase Motions
Before we tackle the general proof, we briefly recall pre-optimal trajectories within the SPM framework while climbing from $N=1$ to $N=3$.
@@ 3.1 Acceleration-Controlled Motions
Let $v$ denote the maximum velocity reached during the motion [^7] Note again, that achieved maximum of the velocity does not need to be equal to the given bound $w_{1}=v_{\max}$; we always have $v\leq w_{1}$., i.e. $v=\max\limits_{t\in[0,T]}|\dot{x}(t)|$. In the triangular case we have
@@ 3.2 Jerk-Controlled Motions
By including the jerk constraint $(N=3)$ we now reach the actual case of the SPM. However, there are four cases to be distinguished (see Figure 2 and Section 5.2). We derive Equation (1) for each of the cases. Let the pre-optimal constant $c$ be denoted by $j$, let $v$ be defined as before in Section 3.1, and let $a:=\max\limits_{t\in[0,T]}|\ddot{x}(t)|$ denote the maximal acceleration reached during the motion.
@@
4-Phase Case. Clearly,
@@
$x(t)$ $=\frac{1}{6}jt_{1}^{3}+\frac{1}{2}v(t-t_{1})+\frac{1}{2}a(t-t_{1})^{2}-\frac{1}{6}j(t-t_{1})^{3}$ (11)
@@
5-Phase Case. Again Equation (7) is valid, but now we have with Equation (11) that
@@
6-Phase Case. In this case,
@@
and via Equation (15) to $\frac{s}{v}=\frac{T}{2}$. Applying Equation (23) leads to $\frac{v}{a}+\frac{a}{j}=\frac{T}{2}$ as well, and hence
$T=\frac{T}{2}+\frac{T}{2}=\frac{s}{v}+\frac{v}{a}+\frac{a}{j}$ (25)
which again proves Equation (1).
7-Phase Case. From the 6-phase case we can already adapt
@@
which completes the proof of Equation (1) for all cases of SPM.
@@ 4 Time Horizon of Pre-Optimal Motions
However, since scanning larger decision trees seems quite unsatisfactory, this section presents a systematic approach for arbitrary orders.
@@
With the notation from Theorem 3, we introduce the earliest time point of achievement as
$T_{n}=\min\left\{t\in[0,T]\penalty 10000\ |\penalty 10000\ x^{(n)}(t)=x_{n}\right\}$ (32)
@@
$T_{n}$ $=$ $\min\left\{t\in[0,T]\penalty 10000\ |\penalty 10000\ x^{(n)}(t)=x_{n}\right\}$ (43)
$=$ $\min\left\{t\in[0,T]\penalty 10000\ |\penalty 10000\ f_{N-n}(t)=f_{N-n}(t_{N-n-1})\right\}$
@@ 5 Applications and Examples
Equipped with the general Equation (33) of Theorem 5 we are able to gain some new insights into of optimal motion planning. In the first subsection we will briefly discuss the cases $N=1$ (velocity-controlled motions) and $N=2$ (acceleration-controlled motions) before the might and elegance of Equation (1) are fully exploited while treating jerk-controlled and snap-controlled movements (Subsections 5.2 and 5.3, respectively). Section 5 closes with remarks on arbitrary order (Subsection 5.4) which culminate in the algorithm for calculating the minimum-time motion (i.e. MTT, Subsection 5.5).
@@ 5.1 Velocity-Controlled and Acceleration-Controlled Motions
  • $\hat{v}_{\mathrm{MTT}}=\hat{v}^{\star}$, $\hat{T}_{\mathrm{MTT}}=\hat{T}^{\star}$, if $\hat{v}^{\star}\leq\hat{v}_{\max}$ (triangular case in Section 3.1, see Figure 1 left), or
  • $\hat{v}_{\mathrm{MTT}}=\hat{v}_{\max}$, $\hat{T}_{\mathrm{MTT}}=\frac{1}{\hat{v}_{\max}}+\hat{v}_{\max}$, otherwise (trapezoidal case, see Figure 1 right).
Figure 3: $\hat{T}$ as a function of $\hat{v}$ (black curve) according to Equation (49) which exhibits a global minimum at $\hat{v}^{\star}=1$ and $\hat{T}^{\star}=2$ (blue cross). The triangular case of MTT occurs if $\hat{v}_{\max}$ lies beyond this minimum (green dashed line). Then the minimal time is exactly $\hat{T}^{\star}$. Otherwise (magenta dashed line, trapezoidal case) the optimum of MTT is located at the magenta bullet. The MDT optimization is colored in orange. The optimum is indicated by the orange bullet.
@@ 5.2 Jerk-Controlled Motions
The “diving fish diagram” of Figure 4 displays $\hat{T}$ as a function of $\hat{v}$ and $\hat{a}$ as a contour plot. There are two red curves – $\hat{a}_{\min}$ curve and $\hat{v}_{\min}$ curve – which delimit the region of all principally possible motions. These curves are defined by the conditions (see Equation (8)):
@@
If such a condition is satisfied the associated order has no cruising phase. For example, $\hat{T}_{1}=2\hat{T}_{2}$ means that $\ddot{x}$ has a triangular shape if it is not equal to 0 (see Figure 2 both upper panels). If there is no cruising phase in $\ddot{x}$, the time for reaching $\hat{v}$ (equal to $\hat{T}_{1}$) is exactly twice the time needed for reaching $\hat{a}$ (equal to $\hat{T}_{2}$). This corresponds to the fact that for every acceleration a certain velocity is quasi automatically achieved. Therefore, not all of the ($\hat{v},\hat{a}$)-pairs are actually possible. And it is precisely these borderline cases that are reflected in the conditions (59, 60).
Furthermore, the global minimum is determined by simultaneously satisfying the equations (59, 60). It is located at $\hat{a}^{\star}=\sqrt[3]{\frac{1}{2}}$, $\hat{v}^{\star}=\sqrt[3]{\frac{1}{4}}$ and $\hat{T}^{\star}=\sqrt[3]{32}$.
Let us now take a closer look at MTT. The four possible cases discussed in Section 3.2 can be easily recovered and reinterpreted within the diving fish diagram. As above, $\hat{v}_{\max}$ and $\hat{a}_{\max}$ are horizontal and vertical cut-offs, respectively. Then:
@@
A slightly different perspective is exhibited in Figure 5, where $\hat{T}$ is displayed as a function of $\hat{T}_{1}$ and $\hat{T}_{2}$ (see Equations (57, 58)). The $\hat{v}_{\min}$ curve has become a straight line. While $\hat{T}_{2}$ is always bounded, $\hat{T}_{1}$ can become arbitrary high. There is a similar procedure for MDT. Since $\hat{T}_{\mathrm{MDT}}$ is given we work on a fixed contour curve.
Figure 5: Dimensionless time horizon $\hat{T}$ as a function of $\hat{T}_{1}$ (dimensionless time for achieving $\hat{v}$) and $\hat{T}_{2}$ (respective time for achieving $\hat{a}$) which is basically a transformation of the diving fish diagram displayed in Figure 4. Note that $\hat{T}=((\hat{T}_{1}-\hat{T}_{2})\hat{T}_{2})^{-1}+\hat{T}_{1}$. The red straight line indicated by $\hat{T}_{1}=2\hat{T}_{2}$ corresponds to the $\hat{v}_{\min}$ curve (see Equation (60)), and the red curve indicated by $\hat{T}_{0}=2\hat{T}_{1}$ belongs to the $\hat{a}_{\min}$ curve (see Equation (59)). This curve is described by $\hat{T}_{1}=\frac{1}{2}\hat{T}_{2}+\sqrt{\frac{1}{4}\hat{T}_{2}^{2}+\hat{T}_{2}^{-1}}$.
@@
All in all, the relevant optimal values can be obtained by targeted solving of algebraic equations. In addition, the diving fish diagram provides an indication of which case actually occurs, i.e. of how many phases the motion consists and which particular bounds are active. Nevertheless, independent of the cases and MTT or MDT, Equation (52) remains overall valid.
@@
Figure 7: Schematic illustration of the solution of MDT with $M=1$, i.e. minimum-velocity motions. In any case, the solution is indicated by an orange bullet. Left: if $\hat{a}_{\max}>\hat{a}^{\star}$, the minimal velocity is achieved at the point where the contour curve belonging to the given time horizon $\hat{T}_{\mathrm{MDT}}$ leaves the admissible region and intersects the $\hat{v}_{\max}$ curve. Right: For small enough values of $\hat{a}_{\max}$ the contour curve reaches the acceleration boundary before the $\hat{v}_{\min}$ curve. The solution is determined by this intersection.